OpenStudy (acespeedfighter):

Zoya asked the students of her class their baseball scores and recorded the scores in the table shown below: Baseball Scores Score Number of Students 0 1 1 2 2 5 3 3 4 6 5 7 6 9 Based on the table, what is the mean baseball score? (1 point) 2.5 2.9 3.5 4.1

OpenStudy (acespeedfighter):

Slender Triple and even quad check your work PLEASE PLEASE PLEASE

OpenStudy (slenderman):

\[0+1+1+2+2+2+2+2+3+3+3+4+4+4+4+4+4+5+5+5+5+5+5+5+6+6+6+6+6+6+6+6+6= X\] x/total is mean

OpenStudy (acespeedfighter):

134

OpenStudy (slenderman):

divided by tthe total is?

OpenStudy (kgrendel0324):

134/the total number of students= mean there are 33 people so 134/33=?

OpenStudy (acespeedfighter):

divide by what?

OpenStudy (kgrendel0324):

33

OpenStudy (kgrendel0324):

134/33=

OpenStudy (kgrendel0324):

make sure you round

OpenStudy (acespeedfighter):

4.1

OpenStudy (acespeedfighter):

D :DD

OpenStudy (acespeedfighter):

that means

OpenStudy (kgrendel0324):

Great :D

OpenStudy (acespeedfighter):

(NUT) (NUT) (NUT) (NUT) (NUT)

OpenStudy (slenderman):

*facepalms self into unconsciousness*

OpenStudy (kgrendel0324):

lol